OpenStudy (anonymous):

I need help with this problem. As a salesperson at Trading Cards Unlimited, Justin receives a monthly base pay plus commission on all that he sells. If he sells $400 worth of merchandise in one month, he is paid $500. If he sells $700 of merchandise in one month, he is paid $575. Find Justin's salary when he sells $2,500 worth of merchandise. Help ASAP, pleeeeaaasee ?

OpenStudy (anonymous):

@alyygirl can you help me please?

OpenStudy (anonymous):

Let b = base pay Let c = commission rate Since you have two unknowns, you need two equations to solve for base pay and commission rate. So the two equations are this: b + 400c = 500 b + 700c = 575 Using the elimination technique multiply the first equation by a negative 1. So the two equations look like this: -b - 400c = -500 b + 700c = 575 Adding the two equations and we this: 300c = 75 c= 75/300 c= .25 So commission rate is .25 or 25% To find the base pay amount, substitute .25 into one of the equations. Which looks like this: b + 700(.25) = 575 b+ 175 = 575 b= 575-175 b= 400 So the base pay is $400 a month. Knowing the base pay and the commission rate we can now calculate how much is earned per month on sales of $2,500. So we have this: 400 + (2,500)(.25) = 1,025 So the paycheck for that month will be $1,025